Jamal's age is 5 years more than 2/5 Karen's age jamal is 17 how old is Karen

Answer by Cromlix(4381)

Hi there,
Make Karen's age = 'x'
5 + 2/5x = 17
2/5x = 17 - 5
2/5x = 12
2x = 5 x 12
2x = 60
x = 60/2
x = 30
Karen is 30 years old.
Hope this helps :-)